### v1

# Euclid Quick Data Release (Q1)

11institutetext: INAF-Osservatorio Astronomico di Capodimonte, Via Moiariello 16, 80131 Napoli, Italy 22institutetext: Department of Physics ”E. Pancini”, University Federico II, Via Cinthia 6, 80126, Napoli, Italy 33institutetext: INFN section of Naples, Via Cinthia 6, 80126, Napoli, Italy 44institutetext: Dipartimento di Fisica e Astronomia ”Augusto Righi” - Alma Mater Studiorum Università di Bologna, via Piero Gobetti 93/2, 40129 Bologna, Italy 55institutetext: INAF-Osservatorio di Astrofisica e Scienza dello Spazio di Bologna, Via Piero Gobetti 93/3, 40129 Bologna, Italy 66institutetext: School of Mathematics, Statistics and Physics, Newcastle University, Herschel Building, Newcastle-upon-Tyne, NE1 7RU, UK 77institutetext: INFN-Sezione di Bologna, Viale Berti Pichat 6/2, 40127 Bologna, Italy 88institutetext: CEA Saclay, DFR/IRFU, Service d’Astrophysique, Bat. 709, 91191 Gif-sur-Yvette, France 99institutetext: Dipartimento di Fisica e Astronomia, Università di Bologna, Via Gobetti 93/2, 40129 Bologna, Italy 1010institutetext: Aix-Marseille Université, CNRS, CNES, LAM, Marseille, France $1$institutetext: Institut d’Astrophysique de Paris, UMR 7095, CNRS, and Sorbonne Université, 98 bis boulevard Arago, 75014 Paris, France 1212institutetext: School of Physics and Astronomy, Sun Yat-sen University, Guangzhou 519082, Zhuhai Campus, China 1313institutetext: Institude for Astrophysics, School of Physics, Zhengzhou University, Zhengzhou, 450001, China 1414institutetext: Institute of Physics, Laboratory of Astrophysics, Ecole Polytechnique Fédérale de Lausanne (EPFL), Observatoire de Sauverny, 1290 Versoix, Switzerland 1515institutetext: SCITAS, Ecole Polytechnique Fédérale de Lausanne (EPFL), 1015 Lausanne, Switzerland 1616institutetext: Dipartimento di Fisica ”E. Pancini”, Universita degli Studi di Napoli Federico II, Via Cinthia 6, 80126, Napoli, Italy 1717institutetext: Institut de Ciències del Cosmos (ICCUB), Universitat de Barcelona (IEEC-UB), Martí i Franquès 1, 08028 Barcelona, Spain 1818institutetext: Institució Catalana de Recerca i Estudis Avançats (ICREA), Passeig de Lluís Companys 23, 08010 Barcelona, Spain 1919institutetext: David A. Dunlap Department of Astronomy & Astrophysics, University of Toronto, 50 St George Street, Toronto, Ontario M5S 3H4, Canada 2020institutetext: Jodrell Bank Centre for Astrophysics, Department of Physics and Astronomy, University of Manchester, Oxford Road, Manchester M13 9PL, UK 2121institutetext: School of Physical Sciences, The Open University, Milton Keynes, MK7 6AA, UK $2$institutetext: Department of Physics and Astronomy, University of British Columbia, Vancouver, BC V6T 1Z1, Canada 2323institutetext: Laboratoire d’etude de l’Univers et des phenomenes eXtremes, Observatoire de Paris, Université PSL, Sorbonne Université, CNRS, 92190 Meudon, France 2424institutetext: Université Paris-Saclay, CNRS, Institut d’astrophysique spatiale, 91405, Orsay, France 2525institutetext: ESAC/ESA, Camino Bajo del Castillo, s/n., Urb. Villafranca del Castillo, 28692 Villanueva de la Cañada, Madrid, Spain 2626institutetext: School of Mathematics and Physics, University of Surrey, Guildford, Surrey, GU2 7XH, UK 2727institutetext: INAF-Osservatorio Astronomico di Brera, Via Brera 28, 20122 Milano, Italy 2828institutetext: Université Paris-Saclay, Université Paris Cité, CEA, CNRS, AIM, 91191, Gif-sur-Yvette, France 2929institutetext: IFPU, Institute for Fundamental Physics of the Universe, via Beirut 2, 34151 Trieste, Italy 3030institutetext: INAF-Osservatorio Astronomico di Trieste, Via G. B. Tiepolo 11, 34143 Trieste, Italy 3131institutetext: INFN, Sezione di Trieste, Via Valerio 2, 34127 Trieste TS, Italy 3232institutetext: SISSA, International School for Advanced Studies, Via Bonomea 265, 34136 Trieste TS, Italy $3$institutetext: Dipartimento di Fisica, Università di Genova, Via Dodecaneso 33, 16146, Genova, Italy 3434institutetext: INFN-Sezione di Genova, Via Dodecaneso 33, 16146, Genova, Italy 3535institutetext: Instituto de Astrofísica e Ciências do Espaço, Universidade do Porto, CAUP, Rua das Estrelas, PT4150-762 Porto, Portugal 3636institutetext: Faculdade de Ciências da Universidade do Porto, Rua do Campo de Alegre, 4150-007 Porto, Portugal 3737institutetext: Dipartimento di Fisica, Università degli Studi di Torino, Via P. Giuria 1, 10125 Torino, Italy 3838institutetext: INFN-Sezione di Torino, Via P. Giuria 1, 10125 Torino, Italy 3939institutetext: INAF-Osservatorio Astrofisico di Torino, Via Osservatorio 20, 10025 Pino Torinese (TO), Italy 4040institutetext: European Space Agency/ESTEC, Keplerlaan 1, 2201 AZ Noordwijk, The Netherlands 4141institutetext: Institute Lorentz, Leiden University, Niels Bohrweg 2, 2333 CA Leiden, The Netherlands 4242institutetext: Leiden Observatory, Leiden University, Einsteinweg 55, 2333 CC Leiden, The Netherlands 4343institutetext: INAF-IASF Milano, Via Alfonso Corti 12, 20133 Milano, Italy $4$institutetext: INAF-Osservatorio Astronomico di Roma, Via Frascati 33, 00078 Monteporzio Catone, Italy 4545institutetext: INFN-Sezione di Roma, Piazzale Aldo Moro, 2 - c/o Dipartimento di Fisica, Edificio G. Marconi, 00185 Roma, Italy 4646institutetext: Centro de Investigaciones Energéticas, Medioambientales y Tecnológicas (CIEMAT), Avenida Complutense 40, 28040 Madrid, Spain 4747institutetext: Port d’Informació Científica, Campus UAB, C. Albareda s/n, 08193 Bellaterra (Barcelona), Spain 4848institutetext: Institute for Theoretical Particle Physics and Cosmology (TTK), RWTH Aachen University, 52056 Aachen, Germany 4949institutetext: Institute for Astronomy, University of Hawaii, 2680 Woodlawn Drive, Honolulu, HI 96822, USA 5050institutetext: Dipartimento di Fisica e Astronomia ”Augusto Righi” - Alma Mater Studiorum Università di Bologna, Viale Berti Pichat 6/2, 40127 Bologna, Italy 5151institutetext: Instituto de Astrofísica de Canarias, Vía Láctea, 38205 La Laguna, Tenerife, Spain 5252institutetext: Institute for Astronomy, University of Edinburgh, Royal Observatory, Blackford Hill, Edinburgh EH9 3HJ, UK 5353institutetext: European Space Agency/ESRIN, Largo Galileo Galilei 1, 00044 Frascati, Roma, Italy 5454institutetext: Université Claude Bernard Lyon 1, CNRS/IN2P3, IP2I Lyon, UMR 5822, Villeurbanne, F-69100, France $5$institutetext: UCB Lyon 1, CNRS/IN2P3, IUF, IP2I Lyon, 4 rue Enrico Fermi, 69622 Villeurbanne, France 5656institutetext: Mullard Space Science Laboratory, University College London, Holmbury St Mary, Dorking, Surrey RH5 6NT, UK 5757institutetext: Departamento de Física, Faculdade de Ciências, Universidade de Lisboa, Edifício C8, Campo Grande, PT1749-016 Lisboa, Portugal 5858institutetext: Instituto de Astrofísica e Ciências do Espaço, Faculdade de Ciências, Universidade de Lisboa, Campo Grande, 1749-016 Lisboa, Portugal 5959institutetext: Department of Astronomy, University of Geneva, ch. d’Ecogia 16, 1290 Versoix, Switzerland 6060institutetext: INAF-Istituto di Astrofisica e Planetologia Spaziali, via del Fosso del Cavaliere, 100, 00100 Roma, Italy 6161institutetext: INFN-Padova, Via Marzolo 8, 35131 Padova, Italy 6262institutetext: Aix-Marseille Université, CNRS/IN2P3, CPPM, Marseille, France 6363institutetext: Space Science Data Center, Italian Space Agency, via del Politecnico snc, 00133 Roma, Italy 6464institutetext: INFN-Bologna, Via Irnerio 46, 40126 Bologna, Italy 6565institutetext: School of Physics, HH Wills Physics Laboratory, University of Bristol, Tyndall Avenue, Bristol, BS8 1TL, UK $6$institutetext: Universitäts-Sternwarte München, Fakultät für Physik, Ludwig-Maximilians-Universität München, Scheinerstrasse 1, 81679 München, Germany 6767institutetext: Max Planck Institute for Extraterrestrial Physics, Giessenbachstr. 1, 85748 Garching, Germany 6868institutetext: INAF-Osservatorio Astronomico di Padova, Via dell’Osservatorio 5, 35122 Padova, Italy 6969institutetext: Institute of Theoretical Astrophysics, University of Oslo, P.O. Box 1029 Blindern, 0315 Oslo, Norway 7070institutetext: Jet Propulsion Laboratory, California Institute of Technology, 4800 Oak Grove Drive, Pasadena, CA, 91109, USA 7171institutetext: Department of Physics, Lancaster University, Lancaster, LA1 4YB, UK 7272institutetext: Felix Hormuth Engineering, Goethestr. 17, 69181 Leimen, Germany 7373institutetext: Technical University of Denmark, Elektrovej 327, 2800 Kgs. Lyngby, Denmark 7474institutetext: Cosmic Dawn Center (DAWN), Denmark 7575institutetext: Max-Planck-Institut für Astronomie, Königstuhl 17, 69117 Heidelberg, Germany 7676institutetext: NASA Goddard Space Flight Center, Greenbelt, MD 20771, USA $7$institutetext: Department of Physics and Astronomy, University College London, Gower Street, London WC1E 6BT, UK 7878institutetext: Department of Physics and Helsinki Institute of Physics, Gustaf Hällströmin katu 2, 00014 University of Helsinki, Finland 7979institutetext: Université de Genève, Département de Physique Théorique and Centre for Astroparticle Physics, 24 quai Ernest-Ansermet, CH-1211 Genève 4, Switzerland 8080institutetext: Department of Physics, P.O. Box 64, 00014 University of Helsinki, Finland 8181institutetext: Helsinki Institute of Physics, Gustaf Hällströmin katu 2, University of Helsinki, Helsinki, Finland 8282institutetext: Centre de Calcul de l’IN2P3/CNRS, 21 avenue Pierre de Coubertin 69627 Villeurbanne Cedex, France 8383institutetext: SKA Observatory, Jodrell Bank, Lower Withington, Macclesfield, Cheshire SK11 9FT, UK 8484institutetext: Dipartimento di Fisica ”Aldo Pontremoli”, Università degli Studi di Milano, Via Celoria 16, 20133 Milano, Italy 8585institutetext: INFN-Sezione di Milano, Via Celoria 16, 20133 Milano, Italy 8686institutetext: Universität Bonn, Argelander-Institut für Astronomie, Auf dem Hügel 71, 53121 Bonn, Germany 8787institutetext: Department of Physics, Institute for Computational Cosmology, Durham University, South Road, Durham, DH1 3LE, UK $8$institutetext: Université Côte d’Azur, Observatoire de la Côte d’Azur, CNRS, Laboratoire Lagrange, Bd de l’Observatoire, CS 34229, 06304 Nice cedex 4, France 8989institutetext: Université Paris Cité, CNRS, Astroparticule et Cosmologie, 75013 Paris, France 9090institutetext: CNRS-UCB International Research Laboratory, Centre Pierre Binetruy, IRL2007, CPB-IN2P3, Berkeley, USA 9191institutetext: Institut d’Astrophysique de Paris, 98bis Boulevard Arago, 75014, Paris, France 9292institutetext: Aurora Technology for European Space Agency (ESA), Camino bajo del Castillo, s/n, Urbanizacion Villafranca del Castillo, Villanueva de la Cañada, 28692 Madrid, Spain 9393institutetext: Institut de Física d’Altes Energies (IFAE), The Barcelona Institute of Science and Technology, Campus UAB, 08193 Bellaterra (Barcelona), Spain 9494institutetext: DARK, Niels Bohr Institute, University of Copenhagen, Jagtvej 155, 2200 Copenhagen, Denmark 9595institutetext: Centre National d’Etudes Spatiales – Centre spatial de Toulouse, 18 avenue Edouard Belin, 31401 Toulouse Cedex 9, France 9696institutetext: Institute of Space Science, Str. Atomistilor, nr. 409 Măgurele, Ilfov, 077125, Romania 9797institutetext: Consejo Superior de Investigaciones Cientificas, Calle Serrano 117, 28006 Madrid, Spain 9898institutetext: Universidad de La Laguna, Departamento de Astrofísica, 38206 La Laguna, Tenerife, Spain $9$institutetext: Dipartimento di Fisica e Astronomia ”G. Galilei”, Università di Padova, Via Marzolo 8, 35131 Padova, Italy 100100institutetext: Institut für Theoretische Physik, University of Heidelberg, Philosophenweg 16, 69120 Heidelberg, Germany 101101institutetext: Institut de Recherche en Astrophysique et Planétologie (IRAP), Université de Toulouse, CNRS, UPS, CNES, 14 Av. Edouard Belin, 31400 Toulouse, France 102102institutetext: Université St Joseph; Faculty of Sciences, Beirut, Lebanon 103103institutetext: Departamento de Física, FCFM, Universidad de Chile, Blanco Encalada 2008, Santiago, Chile 104104institutetext: Universität Innsbruck, Institut für Astro- und Teilchenphysik, Technikerstr. 25/8, 6020 Innsbruck, Austria 105105institutetext: Institut d’Estudis Espacials de Catalunya (IEEC), Edifici RDIT, Campus UPC, 08860 Castelldefels, Barcelona, Spain 106106institutetext: Satlantis, University Science Park, Sede Bld 48940, Leioa-Bilbao, Spain 107107institutetext: Institute of Space Sciences (ICE, CSIC), Campus UAB, Carrer de Can Magrans, s/n, 08193 Barcelona, Spain 108108institutetext: Instituto de Astrofísica e Ciências do Espaço, Faculdade de Ciências, Universidade de Lisboa, Tapada da Ajuda, 1349-018 Lisboa, Portugal 109109institutetext: Cosmic Dawn Center (DAWN) 110110institutetext: Niels Bohr Institute, University of Copenhagen, Jagtvej 128, 2200 Copenhagen, Denmark 111111institutetext: Universidad Politécnica de Cartagena, Departamento de Electrónica y Tecnología de Computadoras, Plaza del Hospital 1, 30202 Cartagena, Spain 112112institutetext: Infrared Processing and Analysis Center, California Institute of Technology, Pasadena, CA 91125, USA 113113institutetext: Dipartimento di Fisica e Scienze della Terra, Università degli Studi di Ferrara, Via Giuseppe Saragat 1, 44122 Ferrara, Italy 114114institutetext: Istituto Nazionale di Fisica Nucleare, Sezione di Ferrara, Via Giuseppe Saragat 1, 44122 Ferrara, Italy 115115institutetext: INAF, Istituto di Radioastronomia, Via Piero Gobetti 101, 40129 Bologna, Italy 116116institutetext: Astronomical Observatory of the Autonomous Region of the Aosta Valley (OAVdA), Loc. Lignan 39, I-11020, Nus (Aosta Valley), Italy 117117institutetext: Department of Physics, Oxford University, Keble Road, Oxford OX1 3RH, UK 118118institutetext: Instituto de Astrofísica de Canarias (IAC); Departamento de Astrofísica, Universidad de La Laguna (ULL), 38200, La Laguna, Tenerife, Spain 119119institutetext: Université PSL, Observatoire de Paris, Sorbonne Université, CNRS, LERMA, 75014, Paris, France 120120institutetext: Université Paris-Cité, 5 Rue Thomas Mann, 75013, Paris, France 121121institutetext: Department of Mathematics and Physics E. De Giorgi, University of Salento, Via per Arnesano, CP-I93, 73100, Lecce, Italy 122122institutetext: INFN, Sezione di Lecce, Via per Arnesano, CP-193, 73100, Lecce, Italy 123123institutetext: INAF-Sezione di Lecce, c/o Dipartimento Matematica e Fisica, Via per Arnesano, 73100, Lecce, Italy 124124institutetext: ICL, Junia, Université Catholique de Lille, LITL, 59000 Lille, France 125125institutetext: ICSC - Centro Nazionale di Ricerca in High Performance Computing, Big Data e Quantum Computing, Via Magnanelli 2, Bologna, Italy 126126institutetext: Instituto de Física Teórica UAM-CSIC, Campus de Cantoblanco, 28049 Madrid, Spain 127127institutetext: CERCA/ISO, Department of Physics, Case Western Reserve University, 10900 Euclid Avenue, Cleveland, OH 44106, USA 128128institutetext: Technical University of Munich, TUM School of Natural Sciences, Physics Department, James-Franck-Str. 1, 85748 Garching, Germany 129129institutetext: Max-Planck-Institut für Astrophysik, Karl-Schwarzschild-Str. 1, 85748 Garching, Germany 130130institutetext: Laboratoire Univers et Théorie, Observatoire de Paris, Université PSL, Université Paris Cité, CNRS, 92190 Meudon, France 131131institutetext: Departamento de Física Fundamental. Universidad de Salamanca. Plaza de la Merced s/n. 37008 Salamanca, Spain 132132institutetext: Université de Strasbourg, CNRS, Observatoire astronomique de Strasbourg, UMR 7550, 67000 Strasbourg, France 133133institutetext: Center for Data-Driven Discovery, Kavli IPMU (WPI), UTIAS, The University of Tokyo, Kashiwa, Chiba 277-8583, Japan 134134institutetext: Dipartimento di Fisica - Sezione di Astronomia, Università di Trieste, Via Tiepolo 11, 34131 Trieste, Italy 135135institutetext: California Institute of Technology, 1200 E California Blvd, Pasadena, CA 91125, USA 136136institutetext: Department of Physics & Astronomy, University of California Irvine, Irvine CA 92697, USA 137137institutetext: Kapteyn Astronomical Institute, University of Groningen, PO Box 800, 9700 AV Groningen, The Netherlands 138138institutetext: Departamento Física Aplicada, Universidad Politécnica de Cartagena, Campus Muralla del Mar, 30202 Cartagena, Murcia, Spain 139139institutetext: Instituto de Física de Cantabria, Edificio Juan Jordá, Avenida de los Castros, 39005 Santander, Spain 140140institutetext: Observatorio Nacional, Rua General Jose Cristino, 77-Bairro Imperial de Sao Cristovao, Rio de Janeiro, 20921-400, Brazil 141141institutetext: Institute of Cosmology and Gravitation, University of Portsmouth, Portsmouth PO1 3FX, UK 142142institutetext: Department of Computer Science, Aalto University, PO Box 15400, Espoo, FI-00 076, Finland 143143institutetext: Instituto de Astrofísica de Canarias, c/ Via Lactea s/n, La Laguna 38200, Spain. Departamento de Astrofísica de la Universidad de La Laguna, Avda. Francisco Sanchez, La Laguna, 38200, Spain 144144institutetext: Caltech/IPAC, 1200 E. California Blvd., Pasadena, CA 91125, USA 145145institutetext: Ruhr University Bochum, Faculty of Physics and Astronomy, Astronomical Institute (AIRUB), German Centre for Cosmological Lensing (GCCL), 44780 Bochum, Germany 146146institutetext: Department of Physics and Astronomy, Vesilinnantie 5, 20014 University of Turku, Finland 147147institutetext: Serco for European Space Agency (ESA), Camino bajo del Castillo, s/n, Urbanizacion Villafranca del Castillo, Villanueva de la Cañada, 28692 Madrid, Spain 148148institutetext: ARC Centre of Excellence for Dark Matter Particle Physics, Melbourne, Australia 149149institutetext: Centre for Astrophysics & Supercomputing, Swinburne University of Technology, Hawthorn, Victoria 3122, Australia 150150institutetext: Department of Physics and Astronomy, University of the Western Cape, Bellville, Cape Town, 7535, South Africa 151151institutetext: DAMTP, Centre for Mathematical Sciences, Wilberforce Road, Cambridge CB3 0WA, UK 152152institutetext: Kavli Institute for Cosmology Cambridge, Madingley Road, Cambridge, CB3 0HA, UK 153153institutetext: Department of Astrophysics, University of Zurich, Winterthurerstrasse 190, 8057 Zurich, Switzerland 154154institutetext: Department of Physics, Centre for Extragalactic Astronomy, Durham University, South Road, Durham, DH1 3LE, UK 155155institutetext: IRFU, CEA, Université Paris-Saclay 91191 Gif-sur-Yvette Cedex, France 156156institutetext: Oskar Klein Centre for Cosmoparticle Physics, Department of Physics, Stockholm University, Stockholm, SE-106 91, Sweden 157157institutetext: Astrophysics Group, Blackett Laboratory, Imperial College London, London SW7 2AZ, UK 158158institutetext: Univ. Grenoble Alpes, CNRS, Grenoble INP, LPSC-IN2P3, 53, Avenue des Martyrs, 38000, Grenoble, France 159159institutetext: INAF-Osservatorio Astrofisico di Arcetri, Largo E. Fermi 5, 50125, Firenze, Italy 160160institutetext: Dipartimento di Fisica, Sapienza Università di Roma, Piazzale Aldo Moro 2, 00185 Roma, Italy 161161institutetext: Centro de Astrofísica da Universidade do Porto, Rua das Estrelas, 4150-762 Porto, Portugal 162162institutetext: HE Space for European Space Agency (ESA), Camino bajo del Castillo, s/n, Urbanizacion Villafranca del Castillo, Villanueva de la Cañada, 28692 Madrid, Spain 163163institutetext: Department of Astrophysical Sciences, Peyton Hall, Princeton University, Princeton, NJ 08544, USA 164164institutetext: INAF-Osservatorio Astronomico di Brera, Via Brera 28, 20122 Milano, Italy, and INFN-Sezione di Genova, Via Dodecaneso 33, 16146, Genova, Italy 165165institutetext: Theoretical astrophysics, Department of Physics and Astronomy, Uppsala University, Box 515, 751 20 Uppsala, Sweden 166166institutetext: Mathematical Institute, University of Leiden, Einsteinweg 55, 2333 CA Leiden, The Netherlands 167167institutetext: Institute of Astronomy, University of Cambridge, Madingley Road, Cambridge CB3 0HA, UK 168168institutetext: Space physics and astronomy research unit, University of Oulu, Pentti Kaiteran katu 1, FI-90014 Oulu, Finland 169169institutetext: Department of Physics and Astronomy, Lehman College of the CUNY, Bronx, NY 10468, USA 170170institutetext: American Museum of Natural History, Department of Astrophysics, New York, NY 10024, USA 171171institutetext: Center for Computational Astrophysics, Flatiron Institute, 162 5th Avenue, 10010, New York, NY, USA
Euclid Quick Data Release (Q1)
LEMON – Lens Modelling with Neural networks. Automated and fast modelling of Euclid gravitational lenses with a singular isothermal ellipsoid mass profile
Euclid Collaboration: V. Busillo valerio.busillo@inaf.it Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    C. Tortora Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    R. B. Metcalf Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    J. W. Nightingale Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    M. Meneghetti Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    F. Gentile Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    R. Gavazzi Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    F. Zhong Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    R. Li Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    B. Clément Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    G. Covone Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    N. R. Napolitano Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    F. Courbin Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    M. Walmsley Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    E. Jullo Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    J. Pearson Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    D. Scott Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    A. M. C. Le Brun Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    L. Leuzzi Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    N. Aghanim Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    B. Altieri Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    A. Amara Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    S. Andreon Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    H. Aussel Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    C. Baccigalupi Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    M. Baldi Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    S. Bardelli Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    P. Battaglia Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    A. Biviano Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    E. Branchini Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    M. Brescia Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    J. Brinchmann Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    S. Camera Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    G. Cañas-Herrera Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    V. Capobianco Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    C. Carbone Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    V. F. Cardone Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    J. Carretero Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    S. Casas Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    M. Castellano Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    G. Castignani Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    S. Cavuoti Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    K. C. Chambers Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    A. Cimatti Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    C. Colodro-Conde Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    G. Congedo Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    C. J. Conselice Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    L. Conversi Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    Y. Copin Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    H. M. Courtois Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    M. Cropper Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    A. Da Silva Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    H. Degaudenzi Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    S. de la Torre Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    G. De Lucia Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    A. M. Di Giorgio Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    J. Dinis Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    H. Dole Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    F. Dubath Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    X. Dupac Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    S. Dusini Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    S. Escoffier Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    M. Farina Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    R. Farinelli Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    F. Faustini Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    S. Ferriol Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    F. Finelli Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    S. Fotopoulou Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    M. Frailis Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    E. Franceschi Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    S. Galeotta Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    K. George Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    W. Gillard Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    B. Gillis Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    C. Giocoli Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    J. Gracia-Carpio Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    B. R. Granett Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    A. Grazian Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    F. Grupp Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    S. V. H. Haugan Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    W. Holmes Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    I. Hook Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    F. Hormuth Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    A. Hornstrup Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    P. Hudelot Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    K. Jahnke Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    M. Jhabvala Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    B. Joachimi Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    E. Keihänen Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    S. Kermiche Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    A. Kiessling Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    B. Kubik Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    M. Kümmel Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    M. Kunz Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    H. Kurki-Suonio Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    Q. Le Boulc’h Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    S. Ligori Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    P. B. Lilje Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    V. Lindholm Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    I. Lloro Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    G. Mainetti Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    D. Maino Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    E. Maiorano Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    O. Mansutti Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    O. Marggraf Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    K. Markovic Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    M. Martinelli Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    N. Martinet Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    F. Marulli Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    R. Massey Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    S. Maurogordato Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    E. Medinaceli Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    S. Mei Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    Y. Mellier Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    E. Merlin Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    G. Meylan Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    A. Mora Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    M. Moresco Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    L. Moscardini Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    R. Nakajima Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    C. Neissner Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    S.-M. Niemi Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    C. Padilla Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    S. Paltani Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    F. Pasian Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    K. Pedersen Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    V. Pettorino Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    S. Pires Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    G. Polenta Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    M. Poncet Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    L. A. Popa Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    L. Pozzetti Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    F. Raison Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    R. Rebolo Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    A. Renzi Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    J. Rhodes Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    G. Riccio Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    E. Romelli Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    M. Roncarelli Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    R. Saglia Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    Z. Sakr Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    A. G. Sánchez Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    D. Sapone Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    B. Sartoris Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    J. A. Schewtschenko Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    M. Schirmer Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    P. Schneider Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    T. Schrabback Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    A. Secroun Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    E. Sefusatti Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    G. Seidel Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    M. Seiffert Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    S. Serrano Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    P. Simon Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    C. Sirignano Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    G. Sirri Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    G. Smadja Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    L. Stanco Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    J. Steinwagner Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    P. Tallada-Crespí Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    A. N. Taylor Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    I. Tereno Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    S. Toft Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    R. Toledo-Moreo Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    F. Torradeflot Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    I. Tutusaus Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    L. Valenziano Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    J. Valiviita Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    T. Vassallo Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    A. Veropalumbo Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    Y. Wang Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    J. Weller Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    G. Zamorani Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    E. Zucca Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    V. Allevato Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    M. Ballardini Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    M. Bolzonella Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    E. Bozzo Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    C. Burigana Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    R. Cabanac Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    M. Calabrese Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    D. Di Ferdinando Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    J. A. Escartin Vigo Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    L. Gabarra Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    M. Huertas-Company Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    S. Matthew Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    N. Mauri Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    A. A. Nucita Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    A. Pezzotta Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    M. Pöntinen Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    C. Porciani Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    V. Scottez Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    M. Tenti Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    M. Viel Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    M. Wiesmann Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    Y. Akrami Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    S. Alvi Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    I. T. Andika Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    S. Anselmi Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    M. Archidiacono Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    F. Atrio-Barandela Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    D. Bertacca Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    M. Bethermin Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    A. Blanchard Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    L. Blot Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    S. Borgani Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    M. L. Brown Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    S. Bruton Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    A. Calabro Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    B. Camacho Quevedo Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    A. Cappi Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    F. Caro Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    C. S. Carvalho Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    T. Castro Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    F. Cogato Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    S. Conseil Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    S. Contarini Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    A. R. Cooray Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    O. Cucciati Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    F. De Paolis Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    G. Desprez Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    A. Díaz-Sánchez Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    S. Di Domizio Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    J. M. Diego Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    P. Dimauro Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    A. Enia Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    Y. Fang Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    A. G. Ferrari Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    P. G. Ferreira Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    A. Finoguenov Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    A. Franco Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    K. Ganga Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    J. García-Bellido Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    T. Gasparetto Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    V. Gautard Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    E. Gaztanaga Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    F. Giacomini Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    F. Gianotti Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    G. Gozaliasl Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    M. Guidi Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    C. M. Gutierrez Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    A. Hall Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    W. G. Hartley Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    S. Hemmati Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    C. Hernández-Monteagudo Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    H. Hildebrandt Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    J. Hjorth Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    J. J. E. Kajava Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    Y. Kang Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    V. Kansal Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    D. Karagiannis Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    K. Kiiveri Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    C. C. Kirkpatrick Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    S. Kruk Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    M. Lattanzi Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    J. Le Graet Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    L. Legrand Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    M. Lembo Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    F. Lepori Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    G. Leroy Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    J. Lesgourgues Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    T. I. Liaudat Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    S. J. Liu Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    A. Loureiro Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    J. Macias-Perez Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    G. Maggio Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    M. Magliocchetti Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    F. Mannucci Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    R. Maoli Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    J. Martín-Fleitas Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    C. J. A. P. Martins Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    L. Maurin Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    M. Miluzio Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    P. Monaco Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    C. Moretti Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    G. Morgante Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    S. Nadathur Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    K. Naidoo Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    P. Natoli Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    A. Navarro-Alsina Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    S. Nesseris Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    F. Passalacqua Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    K. Paterson Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    L. Patrizii Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    A. Pisani Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    D. Potter Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    S. Quai Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    M. Radovich Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    I. Risso Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    P.-F. Rocci Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    S. Sacquegna Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    M. Sahlén Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    E. Sarpa Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    A. Schneider Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    M. Schultheis Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    D. Sciotti Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    E. Sellentin Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    M. Sereno Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    L. C. Smith Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    J. Stadel Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    K. Tanidis Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    C. Tao Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    G. Testera Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    R. Teyssier Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    S. Tosi Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    A. Troja Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    M. Tucci Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    C. Valieri Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    A. Venhola Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    D. Vergani Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    G. Vernardos Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    G. Verza Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    P. Vielzeuf Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    N. A. Walton Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)
The Euclid mission aims to survey around 14 0001400014\,00014 000 $\deg^{2}$ of extragalactic sky, providing around $10^{5}$ gravitational lens images. Modelling of gravitational lenses is fundamental to estimate the total mass of the lens galaxy, along with its dark matter content. Traditional modelling of gravitational lenses is computationally intensive and requires manual input. In this paper, we use a Bayesian neural network, LEns MOdelling with Neural networks (LEMON), for modelling Euclid gravitational lenses with a singular isothermal ellipsoid mass profile. Our method estimates key lens mass profile parameters, such as the Einstein radius, while also predicting the light parameters of foreground galaxies and their uncertainties. We validate LEMON’s performance on both mock Euclid data sets, real Euclidised lenses observed with Hubble Space Telescope (hereafter HST), and real Euclid lenses found in the Perseus ERO field, demonstrating the ability of LEMON to predict parameters of both simulated and real lenses. Results show promising accuracy and reliability in predicting the Einstein radius, axis ratio, position angle, effective radius, Sérsic index, and lens magnitude for simulated lens galaxies. The application to real data, including the latest Quick Release 1 strong lens candidates, provides encouraging results, particularly for the Einstein radius. We also verified that LEMON has the potential to accelerate traditional modelling methods, by giving to the classical optimiser the LEMON predictions as starting points, resulting in a speed-up of up to 26 times the original time needed to model a sample of gravitational lenses, a result that would be impossible with randomly initialised guesses. This work represents a significant step towards efficient, automated gravitational lens modelling, which is crucial for handling the large data volumes expected from Euclid.
Key Words.:
Gravitational lensing: strong, Methods: data analysis, Galaxies: elliptical and lenticular, cD
1 Introduction
Gravitational strong lensing is a rare astrophysical phenomenon that occurs when a massive object, such as a galaxy, bends the light coming from a background distant object. This effect is a prediction of Einstein’s general relativity, and is an important tool in the study of galaxy evolution, given that a gravitational lens can magnify distant galaxies, allowing for the detailed study of structures in galaxies at high redshift (Coe et al. 2013). Moreover, the distortion of the background object gives information on the properties of the deflector. In particular, when the redshifts of source and lens are known, galaxy-galaxy strong lensing allows us to determine with extreme precision the projected total mass within the Einstein radius of the foreground galaxy, without requiring any modelling a priori. By assuming a model for the deflector mass profile, it is possible to infer more quantities related to the dark matter content of galaxies, such as the total dark matter mass and the dark matter fraction (Koopmans et al. 2006; Tortora et al. 2010; Auger et al. 2010; Sengül & Dvorkin 2022), along with other properties, such as the mass density slope and the initial mass function (Gavazzi et al. 2007; Koopmans et al. 2009; Treu et al. 2010; Sonnenfeld et al. 2013; Shajib et al. 2021); finally, through gravitational lensing, it is also possible to detect substructures of the lens galaxy (Vegetti et al. 2010). Galaxy-galaxy strong lensing is also a powerful probe of the small-scale structure of galaxy clusters (e.g., Meneghetti et al. 2020, 2022, 2023).
The Euclid mission will observe around 14 0001400014\,00014 000 $\deg^{2}$ of extragalactic sky from the Sun-Earth Lagrange point L2, surveying of order $10^{9}$ galaxies (for a description of the Euclid Wide Survey, we refer the reader to Euclid Collaboration: Scaramella et al. 2022; Euclid Collaboration: Mellier et al. 2024, while an overall description of the mission can be found in Laureijs et al. 2011).
Following Euclid’s launch in 2023, ESA and collaborators initiated the Early Release Observations programme (ERO, Euclid Early Release Observations 2024; Cuillandre et al. 2024a), targeting 17 astronomical objects, including galaxy clusters, nearby galaxies, globular clusters, and star-forming regions, with the objective of demonstrating the telescope’s capabilities and providing early scientific insights (Atek et al. 2024; Cuillandre et al. 2024b; Hunt et al. 2024; Kluge et al. 2024; Marleau et al. 2024; Massari et al. 2024; Saifollahi et al. 2024). Among these targets, the Perseus cluster field is especially of interest for strong lensing science, with two main projects underway: the ERO Lens Search Experiment (ELSE), where a blind visual search of galaxy-scale strong lensing systems is carried out (Acevedo Barroso et al. 2024), finding 16 lens candidates (of which five have a convincing lens model) and a parallel search of strong lenses via the use of neural networks (Pearce-Casey et al. 2024) aimed at testing their performance on a real Euclid field. A similar search with neural networks in the other ERO fields is carried out in Nagam et al. (2025). The first statistically relevant sample of around $500$ strong lenses in Euclid has finally been collected with the data from the Euclid Quick Data Release 1 (Q1, Euclid Quick Release Q1 2025; Euclid Collaboration: Aussel et al. 2025), which also represents the first publicly available data set of Euclid Wide Survey (EWS)-like images, processed with the same pipeline and reaching the same depth as the final survey. Despite covering only $63.1$ $\deg^{2}$, Q1 provides a powerful testbed for Euclid’s lens searching capabilities, with the strong lensing discovery engine at the forefront of these efforts (Euclid Collaboration: Walmsley et al. 2025; Euclid Collaboration: Rojas et al. 2025; Euclid Collaboration: Lines et al. 2025; Euclid Collaboration: Li et al. 2025; Euclid Collaboration: Holloway et al. 2025).
High-resolution imaging, together with a large field of view, are crucial for detecting a large number of gravitational lenses. However, the substantial amount of data coming from the mission will necessarily require automated methods for both lens finding and lens modelling. Currently, convolutional neural networks (CNNs) are already used extensively in lens-finding tasks, having generally good performance (Petrillo et al. 2017, 2019a, 2019b; Metcalf et al. 2019; Cañameras et al. 2020; Li et al. 2020; Rezaei et al. 2022; Rojas et al. 2022; Nagam et al. 2024). Instead, modelling of gravitational lenses has traditionally relied on complex and time-consuming techniques that require manual inputs, such as maximum likelihood and Markov chain Monte Carlo methods (Keeton 2016; Nightingale et al. 2021a, b; Harvey-Hawes & Wiltshire 2024). The advent of machine learning has allowed for the automation and acceleration of this process. In particular, CNNs and Bayesian neural networks (BNNs) have shown remarkable results in recovering lens parameters (Perreault Levasseur et al. 2017; Pearson et al. 2019, 2021; Schuldt et al. 2021, 2023; Gentile et al. 2023).
The main algorithm for this work is LEns MOdelling with Neural networks (LEMON; Gentile et al. 2023), a BNN able to perform fast automated analysis of strong gravitational lenses. In the original work, Gentile et al. (2023) used LEMON to estimate three parameters of a strong gravitational lens modelled as a singular isothermal ellipsoid (SIE), the Einstein radius and the two components of the ellipticity, for mock HST and Euclid gravitational lenses. Along with the expected values, LEMON is also able to output uncertainties for each parameter.
In this work, we expand the results of Gentile et al. (2023), by predicting both mass and light parameters of the foreground galaxies in isolated galaxy-galaxy gravitational lenses, considering a single Sérsic model for the light profile and using mock Euclid lenses with contaminants in the image for both training and testing. We will also verify the behaviour of both uncertainty components and verify the capability of LEMON to generalise to real Euclidised HST lens images.
The paper is structured as follows. In Sect. 2 we give details on the various data sets used for the analysis. In Sect. 3 we briefly introduce the LEMON algorithm. In Sect. 4 we describe the training procedure, the metrics used to quantify the performance of LEMON, and the calibration procedure for the total uncertainty. Section 5 is dedicated to the results of the analysis on the simulated test set, while Sect. 6 shows the results associated with real Euclidised lenses and real Euclid lenses found in the Perseus ERO and Q1 fields. Section 7 shows how the joint use of LEMON and classical modelling methods could provide a speed up of the latter. We finally give our conclusions in Sect. 8.
2 Data
In the following, we will give details on the various data sets used for the analysis. In Sect. 2.1, the Euclid mock lenses used for training and testing are described. We then describe the real Euclidised lenses used for testing LEMON performance on real lenses in Sect. 2.2.
2.1 Euclid mock lenses
Table 1: List of parameter ranges for the Euclid mock lenses.
Parameters Range
Einstein radius $R_{\textrm{Ein}}/\mathrm{arcsec}$ $[0.51,4.44]$
Axis ratio $q$ $[0.11,1.00]$
[…]
Lens effective radius $R_{\textrm{e, lens}}/\mathrm{arcsec}$ $[0.05,2.00]$
Lens Sérsic index $n_{\mathrm{lens}}$ $[2.00,8.00]$
Lens magnitude $m_{\mathrm{lens}}$ $[13.88,24.68]$
 [^1] Only position angle and Sérsic index are extracted from a uniform distribution. This choice has been made to avoid introducing a prior in the training of the network. All the other parameters are extracted from non-analytic, empirical distributions derived from the Flagship simulation(Euclid Collaboration: Castander et al. 2024). The coverage of the parameters in the reported intervals is thus non-uniform.
Figure 1: Distribution of mass and light parameters for the Euclid mock lenses. Blue bars show the distribution of lenses without contaminants in the image, while orange bars show the distribution of lenses with contaminants.
For training and performance-testing of LEMON, we have made joint use of 50 0005000050\,00050 000 lenses ‘with companions’ (i.e., having contaminants other than the main lens in the image), each of which contributes to the projected lensing potential, and 50 0005000050\,00050 000 lenses ‘without companions’. Both sets are simulated by the Strong Lensing Science Working Group of the Euclid Consortium (Metcalf et al., in prep.). The lens images are made by simulating the Euclid visible instrument (VIS, Euclid Collaboration: Cropper et al. 2024) camera output, at the depth of the Euclid Wide Survey. The two samples are selected to capture the full range of environmental scenarios that may occur when observing a gravitational lens, including objects along the line of sight as well as companion or background objects surrounding the lens system. Separating the two sets allows us to also test whether the lack of training on images with companions affects the recovery of the parameters, given that the presence of companions in the image could be mistaken by the network as part of the lensing system (see App. A).
The lenses are constructed by first considering galaxies in the Euclid Flagship simulation (Euclid Collaboration: Castander et al. 2024), taken from a sky area of the order of $100\,\textrm{deg}^{2}$, under the requirements that the galaxy chosen is brighter than the lens galaxy magnitude limit of 24 in the $I_{\scriptscriptstyle\rm E}$ band, which is the VIS imaging band, and categorised as a central galaxy. Central galaxies are galaxies that are positioned at the centres of the dark matter halos. Large halos also contain satellite galaxies, which should not produce isolated lenses like the ones we wish to model here. For each selected candidate, the apparent magnitude, effective radius, position angle, axis ratio and redshift are taken from the Flagship and converted into a surface brightness model. In general, the light model is composed of a disc and a spheroid component, but in our catalogue, only galaxies with a spheroid component are considered. This spheroid component is represented by a single Sérsic profile. The Sérsic index is uniformly distributed between 2 and 8. A mass model is then chosen by considering analytical profiles centred on each galaxy. In our catalogue, all lens mass profiles are considered to be SIE profiles.
The mass in the lens is normalised with the ratio of dark matter to stellar mass within one effective radius. The stellar mass and effective radius are taken from the Flagship simulation. The dark matter fraction within one effective radius is taken to be normally distributed with a mean of 0.6 and a standard deviation of 0.1 consistent with observations (Mukherjee et al. 2022). The axis ratio and position angle of the mass is set to those of the light.
Once the lens galaxies have been selected, a source galaxy is placed in the relative background. This increases the sample size of lenses in the data set, in order to offset the inherent rarity of ‘natural lenses’ (i.e. lenses in the Flagship simulation that naturally have a source behind them when the image is traced onto the observer plane). The redshift of the source is sampled randomly from the following distribution:
$p(z)\propto z^{\gamma}\textrm{e}^{-z^{2}/z_{0}^{2}}\,,$ (1)
where $\gamma=-0.23$ and $z_{0}=3.06$ are values found by fitting to the COSMOS 2020 Farmer catalogue $I$-band number counts (Weaver et al. 2022). The surface brightness is represented by between one to four Sérsic profiles. To emulate some of the complex morphologies of high-redshift sources, the number of Sérsic profiles, their relative brightnesses, and their positions are generated randomly from distributions tuned by eye. Their total brightnesses and overall sizes are based on randomly selected sources from the Hubble Ultra Deep Field (HUDF) with similar redshifts (see Meneghetti et al. 2008, 2010 for a discussion of this sample).
Ray-shooting through the light cones is performed, to construct an image of the lens. The deflection caused by other objects along the line of sight is not included, so that the mass distribution remains an SIE in all cases. The lens is then rejected or accepted based on an observability criterion. This criterion seeks to filter out lenses that would not be recognisable as lenses. It combines the signal-to-noise ratio (S/N) in the lensed images, the contrast between the source and lens surface brightnesses and the morphology of the lensed source images. More details on the procedure are available in Metcalf et al. (in prep.).
[…]
The six parameters listed in Table 1 are the main ones that we want to predict with LEMON (the recovery of a seventh parameter, the Einstein mass, is discussed in App. B). They are listed as follows.
  1. 1.
    Einstein radius, in units of arcseconds, obtained via the formula
    $R_{\textrm{Ein}}=\frac{$648\,000$}{\pi}\,\sqrt{\frac{A_{\textrm{Ein}}}{\pi}}\,,$ (2)
    where $A_{\textrm{Ein}}$ is the Einstein area in square radians, defined as the area of the largest critical curve, and the factor 648 000648000648\,000648 000 corresponds to converting $\pi$ radians into arcseconds.
  2. 2.
    Axis ratio, defined as $q=b/a$, where $b$ and $a$ are, respectively, the semi-minor and semi-major axes of the ellipsoidal distribution.
  3. 3.
    Position angle, defined as the clockwise angle of the semi-major axis of the ellipsoidal distribution, starting from left to top with respect to the image, in units of degrees.
  4. 4.
    Circularised effective radius of the lens $R_{\textrm{e}}$, i.e., the radius enclosing one-half of the lens galaxy light, obtained from the major axis effective radius $R_{\textrm{e, maj}}$ via the formula $R_{\textrm{e}}=\sqrt{q}\,R_{\textrm{e, maj}}$. We will refer to it as $R_{\textrm{e, lens}}$ in the following.
  5. 5.
    Lens Sérsic index, which regulates the slope of the surface brightness profile. It corresponds to the parameter $n$ that appears in the Sérsic profile definition
    $I(r)=I_{\textrm{e}}\exp\left\{-b_{n}\left[\left(\frac{r}{R_{\textrm{e}}}\right)^{1/n}-1\right]\right\}\,,$ (3)
    where $I_{\textrm{e}}=I(r=R_{\textrm{e}})$ and $b_{n}$ is a parameter, solution to the equation $\Gamma(2n)=2\,\gamma(2n,b_{n})$, where $\Gamma$ and $\gamma$ are the complete and incomplete gamma functions, respectively. We will refer to it as $n_{\textrm{lens}}$ in the following.
  6. 6.
    Lens magnitude, defined as the apparent magnitude in the $I_{\scriptscriptstyle\rm E}$ band of the main lens.
It should be noted that, in these simulations, we implicitly assume that the ellipticity and orientation of the mass profile of the lens follows those of the light profile. This assumption generally holds for early-type galaxies, but outliers can be found in real scenarios, mainly due to the presence of faint discs or external galaxies that induce a substantial external shear (see e.g., Gavazzi et al. 2012). This assumption can thus lead to systematic errors in predicting lens parameters of systems where the mass does not follow the light profile, because the network tends to learn to estimate the parameters from the light distribution, rather than from the lensed image of the source, as noted in Pearson et al. (2019). In future works, we also plan to train on lenses for which this assumption does not hold, as was done in Gentile et al. (2023), albeit with more simplistic mock lenses. The original simulated images have dimensions of $20\arcsec\times 20\arcsec$, as shown in Fig. 2, but during the training of LEMON we create $10\arcsec\times 10\arcsec$ cutouts of these images, centred on the lenses, in order to better see the lens features.
Figure 2: $20\arcsec\times 20\arcsec$ original VIS images of some examples of Euclid mock lenses used for training LEMON. Top row: lenses with companions in the image. Bottom row: lenses without companions.
2.2 Euclidised lenses
In addition to the set of simulated images, we test the performance of LEMON using a sample of known strong gravitational lenses. These lenses are observed during several strong lensing surveys with HST or serendipitously discovered in the HST archive. Section 6.1 gives more details on the data set. We use the code HST2EUCLID  (Bergamini et al., in prep.) to Euclidise the HST images, mimicking Euclid observations of these sources at the same depth and resolution of the EWS in the $I_{\scriptscriptstyle\rm E}$ band.
We summarise here the procedure implemented in HST2EUCLID. For more details, we refer the reader to Bergamini et al. (in prep.). First, we convert the pixel values of the HST images from units of electrons per second (${\rm e}\,{\rm s}^{-1}$) to physical flux densities (${\rm erg}\,{\rm s}^{-1}\,{\rm cm}^{-2}\,{\rm Hz}^{-1}$). We perform this operation for all available HST images in photometric bands overlapping with the $I_{\scriptscriptstyle\rm E}$ passband. For the lenses in the data set used in this paper, we use only HST observations with the Advanced Camera for Surveys (ACS) in the F814W band.
In the second step, we account for the Euclid PSF. Since the input images are already convolved with the HST PSF, we do not convolve them directly with the Euclid PSF model. Instead, we create a convolution kernel that matches the HST to the Euclid PSF. Since we could not measure the PSF for each HST lens, we use a model for the ACS observations in the F814W band computed with the software TinyTim. [^2] https://www.stsci.edu/hst/instrumentation/focus-and-pointing/focus/tiny-tim-hst-psf-modeling The matched kernel for each pair of HST and Euclid PSFs is computed with the Python function create_matching_kernel of the Photutils package [^3] https://photutils.readthedocs.io/en/stable (Bradley et al. 2023). We then re-bin the images from the HST pixel grid to the correct Euclid pixel scale. The resulting images have pixel scales of $100\,{\rm mas}\,{\rm px}^{-1}$ in the $I_{\scriptscriptstyle\rm E}$ band. The re-binned images, expressed in units of physical flux densities, are then converted into units of electrons per second using the Euclid zero point.
Finally, we add Poisson photon noise. To accomplish this task, we assume that the noise in the input HST background subtracted images is negligible compared to the noise in the Euclid observations. This approximation is supported by the significantly greater depth of the HST observations compared with the Euclid images. We adjust the noise level to achieve the nominal S/N of 10 for an extended source of magnitude 24.5 within an aperture of $0.65″$ radius expected for the EWS (Euclid Collaboration: Scaramella et al. 2022).
3 Introduction to LEMON
LEMON (Gentile et al. 2023) is a BNN (Charnock et al. 2020), a machine-learning algorithm that employs the feature-recognition capabilities of a CNN to predict model parameters of a gravitational lens, using Bayesian statistics to estimate the respective uncertainties. The key point of LEMON is its ability, by using BNNs, to estimate the two main sources of uncertainty in machine-learning inference applications, namely epistemic uncertainty and aleatoric uncertainty.
  • Aleatoric uncertainty: also known as statistical uncertainty, is associated with the intrinsic quality of the data analysed by the algorithm. The most common sources of aleatoric uncertainty are a low S/N, corruptions in the images (e.g., masked regions) and source blending. This kind of uncertainty cannot be reduced with training, because it depends on the quality of the image being analysed.
  • Epistemic uncertainty: also known as systematic uncertainty, it is caused by a lack of knowledge about the best model. It refers to the ignorance of the machine-learning algorithm about a certain region of the parameter space, and hence to the state of the machine rather than to the state of the images. This kind of uncertainty can be reduced with a higher completeness of the training set.
The aleatoric uncertainty for a certain image is given as an output by LEMON, together with the predicted parameter value (for details on how the aleatoric uncertainty is predicted by LEMON, refer to Gentile et al. 2023). The epistemic uncertainty for an image can be obtained thanks to the dropout layers of LEMON: by randomly switching off some connections between the neurons of the neural network, predicting repeatedly the parameters associated to a single image is equivalent to sampling from the posterior probability of the parameters. The epistemic uncertainty can then be evaluated by measuring the standard deviation of this sampled distribution.
The total uncertainty for a generic parameter $p$ can be obtained via the following procedure (Gentile et al. 2023). For each image, LEMON predicts the mean value, $\overline{p}$, and the aleatoric uncertainty, $\sigma_{\textrm{A}}$, for $p$. A value $p_{i}$ is then sampled from a Gaussian distribution centred on $\overline{p}$ with a standard deviation equal to $\sigma_{\textrm{A}}$. This is repeated $\mathcal{N}$ times (we used $\mathcal{N}=50$ to optimise for speed), sampling the posterior distribution of the parameter $p$. Finally, we compute the standard deviation of the distribution associated with the extractions of $p_{i}$, obtaining the total uncertainty.
LEMON employs the ResNet architecture (He et al. 2015), a deep residual network known for its success on computer vision tasks (Russakovsky et al. 2015). The architecture of LEMON is composed of: convolutional layers, used for feature extraction; dropout layers, to facilitate Bayesian uncertainty estimation through Monte Carlo sampling; finally, a fully connected layer made by $2N_{\textrm{param}}$ nodes, with $N_{\textrm{param}}$ number of parameters estimated by the network. A corresponding number of aleatoric uncertainties is estimated through the minimisation of a Gaussian log-likelihood loss function, with an additional regularisation term in order to prevent overly large or small variance estimates. Technical details on the implementation can be found in Gentile et al. (2023).
4 Methods
In this section, we will first describe the training procedure for LEMON in Sect. 4.1. Section 4.2 is dedicated to the description of the metrics that we used to evaluate the performance of LEMON. Finally, Sect. 4.3 is dedicated to discussing the calibration procedure for the total uncertainty.
4.1 Training the network
We prepare the training of LEMON by taking the full sample of 100 000100000100\,000100 000 mock Euclid lenses, both with and without companions in the image, and splitting it into a training set formed by 80 0008000080\,00080 000 lenses (equally split between companions and non-companions), a validation set and a test set formed by 10 0001000010\,00010 000 lenses each (equally split as before). We shuffle the three sets to avoid possible biases during learning. Differently from Gentile et al. (2023), we use the ResNet-50 architecture (He et al. 2015) when implementing LEMON. The increased number of layers with respect to ResNet-34 should improve the accuracy in the recovery of the parameters. Before feeding the images into LEMON, we apply a preprocessing step where pixel values are first rescaled to the range [0,1], and then transformed using a square root scaling to enhance the visibility of low-intensity features.
Training is performed via stochastic gradient descent, with a batch size of 64 images. For the optimiser, we used ADAM (Kingma & Ba 2015), with a starting learning rate of $10^{-4}$. To prevent overfitting, we used both the ReduceLROnPlateau and EarlyStopping callbacks of the keras Python library. The former reduces the learning rate of the optimiser when no improvement is seen for a given number of epochs, while the latter stops the training when the validation loss stops decreasing. Dropout layers, needed for estimating the epistemic uncertainty, are employed at both training and testing time, with a drop rate of 0.05.
We performed the training via the tensorflow Python library (Abadi et al. 2015), using a single Nvidia RTX 4080 graphics processing unit (GPU). The whole training procedure required 15 hours and 30 minutes, with a total of 70 epochs.
4.2 Metrics
To evaluate the performance of LEMON in recovering the lens parameters, we make use of various statistical estimators.
  • Bias, defined as
    $\mu(\hat{\mathbf{y}},\mathbf{y}):=\textrm{median}(\hat{\mathbf{y}}-\mathbf{y})\,,$ (4)
    where $\hat{\mathbf{y}}$ and $\mathbf{y}$ are the arrays of the predicted and the true parameter values, respectively. [^4] A single predicted value $\hat{y}_{i}$ is the mean of the distribution of the $\mathcal{N}$ parameter extractions $p_{j}$ described in Sect. 3. This is an indicator of the accuracy of the prediction: higher bias values indicate a systematic overestimation of the predictions, and vice versa for lower bias values.
  • Discrepancy distribution scatter: the scatter of the discrepancy distribution is obtained from the 16th and 84th percentiles of the discrepancy distribution via the following definitions:
    $\sigma^{-}$ $:=\textrm{median}(\hat{\mathbf{y}}-\mathbf{y})-\textrm{percentile}_{\textrm{16th}}(\hat{\mathbf{y}}-\mathbf{y})\,,$ (5)
    $\sigma^{+}$ $:=\textrm{percentile}_{\textrm{84th}}(\hat{\mathbf{y}}-\mathbf{y})-\textrm{median}(\hat{\mathbf{y}}-\mathbf{y})\,.$ (6)
    Bigger values of these metrics indicate higher scatter from the ideal $\hat{y}=y$ relation. Since they are obtained from the percentiles, these quantities are robust estimators of the scatter.
  • Root mean square error (RMSE), defined as
    $\textrm{RMSE}:=\sqrt{\frac{1}{N}\sum_{i=1}^{N}\left(\hat{y_{i}}-y_{i}\right)^{2}}\,,$ (7)
    where $\hat{y}_{i}$ and $y_{i}$ are the generic components of the $\hat{\mathbf{y}}$ and $\mathbf{y}$ arrays, respectively, which have length $N$. This metric is a measure of predictive power: a value of zero indicates a perfect fit to the true parameter values.
  • Mean absolute error (MAE), defined as
    $\textrm{MAE}:=\frac{1}{N}\sum_{i=1}^{N}\left|\hat{y}_{i}-y_{i}\right|\,.$ (8)
    This metric is similar, but not equivalent to RMSE. It is the average absolute vertical distance between each point $(y_{i},\hat{y}_{i})$ and the ideal $\hat{y}=y$ relation.
  • Normalised median absolute deviation (NMAD), defined as
    $\textrm{NMAD}:=1.48\,\textrm{median}[\left|(\hat{\mathbf{y}}-\mathbf{y})-\textrm{median}(\hat{\mathbf{y}}-\mathbf{y})\right|]\,.$ (9)
    Figure 3: Reliability plots for the parameters predicted by LEMON. Left: trends before the calibration procedure. Right: trends after the calibration procedure.
    The NMAD is roughly analogous to the standard deviation ($\sigma\approx\kappa\,\textrm{MAD}$, where MAD is the median absolute deviation and $\kappa$ is a constant scale factor, which is approximately equal to $1.48$ for normally distributed data). An advantage of the NMAD metric is that it is a non-parametric estimator, which is robust to outliers.
  • Coefficient of determination, defined as
    $R^{2}:=1-\frac{\sum_{i=1}^{N}\left(\hat{y}_{i}-y_{i}\right)^{2}}{\sum_{i=1}^{N}\left(y_{i}-\overline{y}\right)^{2}}\,,$ (10)
    where
    $\overline{y}=\frac{1}{N}\sum_{i=1}^{N}y_{i}$ (11)
    is the mean value of the true parameter over all the lenses. This metric provides a measure of how well the parameters are predicted by LEMON. Given a fixed parameter, if it is perfectly predicted by the network, $\hat{y}_{i}=y_{i}$ for all the points of the data set, and thus $R^{2}=1$. A baseline network, which always outputs the mean value of the true parameter (i.e., $\hat{y}_{i}=\overline{y}$), will produce $R^{2}=0$. If the network performs worse than the baseline model (i.e., the mean of the data provides a better fit than than the predicted values), the values of $R^{2}$ will be negative. [^5] Notice that the coefficient of determination is not the square of a quantity $R$, and therefore can also assume negative values.
4.3 Calibration procedure
Uncertainties produced by BNNs are known to be affected by a systematic over- or under-estimation of the confidence intervals (Guo et al. 2017). Therefore, a calibration procedure for the uncertainties is required. Following Gentile et al. (2023), we implement the Platt-scaling method (Kull et al. 2017), which consists in rescaling the predicted total uncertainty of a given parameter by a factor $s:\sigma\rightarrow s\sigma$, in order to match the empirical cumulative distribution function (CDF) with the CDF of a Gaussian distribution.
Figure 4: Recovery plot of the parameters of Euclid mock lenses, showing how well the predicted model parameters from LEMON ($y$ coordinate) reproduce the corresponding true value ($x$ coordinate), for each lens of the test set (blue points). The ideal recovery line is shown as a dashed black line. The median trend of the scatter plots, along with the respective scatters associated with the 16th and 84th percentiles, are shown as red points and bars, respectively. For each panel, we also report the respective cumulative metrics. The bottom two rows of each parameter show the bias and the NMAD, respectively, as defined in Sect. 4.2, as a function of the true value of the parameters estimated by LEMON. For clarity, we have plotted only $2000$ random points from the test set.
To obtain the scaling factors, we first proceed by splitting the validation set into $70\%$ for the calibration training set and $30\%$ for the calibration test set. We then build the uncalibrated reliability plot, by putting on the $x$-axis the values extracted from the CDF of a Gaussian distribution and on the $y$-axis the respective values extracted from the actual CDF. We fit the resulting trend with a $\beta$-function, and find the Platt-scaling factor, $s$, by minimising the difference between the $\beta$-function and the bisector of the reliability plot. This is done for each parameter, obtaining the values reported in Table 2.
Table 2: Rescaling factors used for the calibration procedure of each parameter predicted by LEMON.
Parameter Scaling factor
Einstein radius $R_{\textrm{Ein}}/\textrm{arcsec}$ $0.99$
Axis ratio $q$ $0.97$
[…]
Lens effective radius $R_{\textrm{e, lens}}/\textrm{arcsec}$ $1.01$
Lens Sérsic index $n_{\textrm{lens}}$ $1.07$
Lens magnitude $m_{\textrm{lens}}$ $1.02$
Figure 3 shows the uncalibrated reliability plot for each parameter in the left panel, and the calibrated reliability plot on the right panel. We can see from the figure that the calibration procedure has its biggest effect on the position angle and Sérsic index uncertainties, which were previously systematically overestimated and underestimated, respectively.
5 Results for the test set
In the following, we will present the results of the analysis on the mock Euclid test set. In Sect. 5.1, we show the recovery of the parameters, along with the trends for the bias and the NMAD for each parameter. In Sect. 5.2, we analyse the behaviour of the uncertainties for each parameter. In App. A, we show the degradation in performance if one uses lenses without companions for training to predict parameters of lenses with companions in the cutouts.
5.1 Recovery of parameters for Euclid mock lenses
Figure 5: Random selection of 20 $10\arcsec\times 10\arcsec$ cutouts of simulated lenses taken from the test set. Red circles show the best predictions for the Einstein radius from LEMON, with corresponding 16th and 84th percentiles shown with dashed red circles. Yellow circles show the true values for the Einstein radii. Predictions and true values are very similar, so in some cases the corresponding circles completely overlap.
Figure 4 shows the recovery of the parameters of the Euclid mock lenses for the test set, for both the mass and light profiles, with the bottom two rows of each parameter showing the bias and NMAD metrics in sequential bins. Overall, the recovery of the parameters for the simulated Euclid lenses of the test set is good, as shown for example in Fig. 5, where we compare the Einstein radius predictions for a random selection of 20 test set lenses with the corresponding ground truth.
[…]
5.2 Behaviour of LEMON uncertainties
Figure 6: Calibrated relative uncertainty as a function of the true value of the lens parameters. For the position angle and lens magnitude, we consider the absolute uncertainty. For clarity, we plotted only $2000$ random points. The median trends of the scatter plots, along with the respective scatters associated with the 16th and 84th percentiles, are shown as red points and bars, respectively.
Figure 7: Trends of aleatoric (blue) and epistemic (orange) relative uncertainties as a function of the true value of the lens parameters. For the position angle and lens magnitude, we consider the absolute uncertainties. For clarity, we plotted only $2000$ random points from the test set.
[…]
We also verified whether the uncertainty on the parameters changes as a function of both S/N and number of companions in the test set images. For the S/N, we use the parameter $n_{\textrm{pix, source}}$, which gives the number of pixels in the lensed source image above $4\sigma$. Results show that the uncertainty for the Einstein radius slightly decreases with an increasing value of $n_{\textrm{pix, source}}$, while the uncertainties for all the other parameters slightly increase. There is, instead, no correlation between uncertainty and number of companions in the images.
The behaviour of the aleatoric and epistemic relative uncertainties associated to the test set is shown in Fig. 7, where the estimated uncertainties of both types are plotted for each lens as a function of the true values of the respective parameters. From the figure, we see that the epistemic uncertainties are systematically lower than the corresponding aleatoric uncertainties. This is the expected behaviour: if the training set correctly spans all the parameter space, the epistemic uncertainties should be as close to zero as possible.
6 Results for real lenses
6.1 LEMON applied to Euclidised lenses
We now apply LEMON to a sub-sample of real Euclidised lenses described in Sect. 2.2. This sub-sample is formed by four different main catalogues, from which the ground truth values have been taken (where available).
  • A sample of 29 Sloan Lens Advanced Camera for Surveys (SLACS) catalogue systems, described in Bolton et al. (2008) and Auger et al. (2009).
  • The sample of 13 early-type/early-type lens systems (EELs) presented in Oldham et al. (2017).
  • A sample of five lenses found in the COSMOS survey (Scoville et al. 2007), detailed in Faure et al. (2008).
  • A sample of 13 lenses taken from those found by visually inspecting the whole imaging data taken with the ACS through the F814W filter up to the 31st of August 2011 (Pawase et al. 2014). Notice that the Einstein radius for these lenses is not reported, and as such we used the radius of the arc (in arcseconds) as a substitute.
All of the Euclidised lenses used in this work have been spectroscopically confirmed. The recovery plot for these lenses is shown in Fig. 8, while the cumulative metrics for each parameter are reported in Table 3. A panel with some of the lenses from each catalogue is shown in Fig. 9.
Figure 8: Recovery plot for the parameters of the Euclidised lenses, showing how well the predicted model parameters from LEMON ($y$ coordinate) reproduce the corresponding true value ($x$ coordinate), for each lens. The ideal recovery line is shown as a dashed black line. The SLACS and ACS best values for the Sérsic indexes are evaluated as the median of the predictions against the value for a De Vaucouleurs profile ($n_{\textrm{lens}}=4$), with uncertainty given by the NMAD of the predictions. Lenses in panel (d) with true values bigger than $2\arcsec$ have been removed due to being outside the training space of LEMON.
Figure 9: Random selection of 20 $10\arcsec\times 10\arcsec$ cutouts of Euclidised lenses taken from the SLACS, EEL, COSMOS, and ACS subsamples. Red circles show the best predictions for the Einstein radius from LEMON, with corresponding 16th and 84th percentiles shown with dashed red circles. Yellow circles show the values for the Einstein radii reported in Bolton et al. (2008), Oldham et al. (2017), Faure et al. (2008), and Pawase et al. (2014), respectively. For ACS lenses, the radius of the arc is taken instead, due to lack of Einstein radius measurement.
Figure 10: Distributions of the Einstein radius (left panel) and the lens effective radius (right panel) as a function of lens redshift. Black regions show the training set, while coloured points show the position of the Euclidised lenses.
Table 3: Cumulative metrics associated to the recovery of the parameters of the Euclidised lenses.
[…]
Bias ($\mu$) 0.17 (0.14) 0.02 (0.01) --1.93 (--2.36) 0.19 (0.22) 0.47 (0.96)
Lower discr. scatter ($\sigma^{-}$) 0.19 (0.12) 0.10 (0.08) 11.18 (11.01) 0.51 (0.53) 1.30 (1.48)
Upper discr. scatter ($\sigma^{+}$) 0.39 (0.24) 0.10 (0.12) 12.77 (9.60)0 0.13 (0.11) 2.57 (2.52)
RMSE 0.63 (0.48) 0.11 (0.11) 32.47 (26.51) 0.38 (0.39) 1.94 (2.13)
MAE 0.42 (0.30) 0.09 (0.08) 18.10 (13.55) 0.32 (0.32) 1.46 (1.63)
NMAD 0.23 (0.14) 0.11 (0.11) 09.11 (10.13) 0.23 (0.22) 1.54 (2.06)
Coefficient of determination ($R^{2}$) --0.03 (0.91)0 0.51 (0.76) 0.57 (0.98) 0.49 (0.53) --0.16 (--0.02)
 [^7] Values in parentheses correspond to the values of the cumulative metrics on the subset of lenses with predicted low uncertainty on the Einstein radius ($\sigma_{R_{\textrm{Ein}}}\leq 0.5\arcsec$).
[…]
In order to try and reduce the noise induced on the cumulative metrics by lenses with low S/N, we re-evaluated them by removing all lenses having predicted uncertainty on the Einstein radius bigger than $0.5″$. Results are shown in the parentheses of Table 3. We can see that the metrics of Einstein radius, axis ratio, and position angle improve substantially, with the values of $R^{2}$ reaching values of up to $0.98$ for the latter. There is, however, no improvement on the metrics of the other parameters. When we remove uncertain effective radii or Sérsic indices, though, the statistical indicators for these parameters do not improve significantly.
[…]
We also checked the distribution of the Einstein and effective radii as a function of the lens redshifts, as can be seen in Fig. 10. From these, we can see that the parameters for the Euclidised lenses are often on the edge of the training-set distribution, if not completely outside it. This could influence negatively the recovery of these parameters, given that the network is not trained to recognise this region of the parameter space. In particular, in panel (d) of Fig. 8, we removed lenses with true effective radii larger than $2\arcsec$, since they fall outside the training sample of LEMON, thus making their predicted values unreliable extrapolations. Furthermore, the true estimates of the effective radii cannot always be compared homogeneously with the predictions, as SLACS and ACS lenses are modelled with a fixed De Vaucouleurs profile (i.e., $n_{\textrm{lens}}=4$), while LEMON considers Sérsic profiles with variable values of $n_{\mathrm{lens}}$. Additionally, the fitting process should ideally remove the arc to obtain an accurate estimate; otherwise, the Sérsic fit may overestimate the radius due to bumps in the surface brightness profile. In future works, we will expand the distribution of the training set to cover more of the parameter space, and check if the recovery of these parameters improves.
Figure 11: $10\arcsec\times 10\arcsec$ cutouts of the five Euclid gravitational lenses found in the Perseus ERO field and modelled in Acevedo Barroso et al. (2024). Top row: unedited cutouts centred on the lenses. Bottom row: the same cutouts, with the predicted Einstein radius (red circles, with dashed circles showing the uncertainty bands) and the value obtained from the classical modelling (yellow circle) superimposed on them.
6.2 LEMON applied to real Euclid ERO lenses
We also applied LEMON to real Euclid lenses found in the Perseus ERO field and modelled in Acevedo Barroso et al. (2024). The modelling has been performed with the pronto software (Vegetti & Koopmans 2009; Rybak et al. 2015a, b; Rizzo et al. 2018; Ritondale et al. 2019; Powell et al. 2021), with the lens mass distribution modelled by an SIE and the light distribution modelled as a composite of three Sérsic profiles. The surface brightness distribution of the source is reconstructed using a pixelated, free-form approach. Regularisation is applied to penalise large gradients in the reconstructed source brightness. The parameters for the lens mass and light profiles are optimised nonlinearly using the MultiNest algorithm (Feroz et al. 2009). The positions of assumed lens images are provided as inputs, with models constrained to ensure that these positions align in the source plane with a tolerance of 11″1  ″. Each model is checked against three criteria: the presence of an SIE critical curve consistent with observed lens image; the alignment of the critical curve with the lens galaxy’s light profile; and finally, the consistency of the reconstructed source with a compact object inside a caustic.
From all the lenses found in the field, we considered the ones that have a convincing model (identified as ‘valid’ in the ‘modelling’ column of table 2 in Acevedo Barroso et al. 2024), and for which the value of the Einstein radius is available. An image of these lenses is shown in Fig. 11. The recovery plot associated to these lenses is shown in Fig. 12. From the figure, we can see that the recovery of the Einstein radius is within $1\sigma$ from the ideal equality line for all lenses, except for the lens J031959+414229. Overall, the recovery of the parameter is sufficiently accurate for real Euclid lenses. It is remarkable that the lens J031749+420011 is recovered correctly, albeit with a large uncertainty, due to the fact that there is a very bright star present in the cutout, which increases both the aleatoric uncertainty and the epistemic uncertainty, the former due to low visibility of the lens and the latter due to not having trained LEMON with stars as contaminants in the image.
Figure 12: Same as Fig. 8, but for the Euclid lenses found in the Perseus ERO field.
6.3 LEMON applied to real Euclid Q1 lenses
We have finally applied LEMON to the sample of strong lens images found by the strong lensing discovery engine in the Euclid Quick Release 1 (Q1) data (Euclid Collaboration: Walmsley et al. 2025). Through an initial scan of the Q1 galaxy catalogues with machine-learning models, followed by citizen science inspection, expert vetting, and system-by-system modelling, Euclid Collaboration: Walmsley et al. (2025) has collected 500 strong lens candidates, divided in Grade A (246) and B (254) lens candidates after the expert vetting scoring. We complement these data with 38 Grade A and 40 Grade B candidates found after the expert visual classification performed on high-velocity dispersion galaxies in Euclid Collaboration: Rojas et al. (2025).
The strong lenses in Euclid Collaboration: Walmsley et al. (2025); Euclid Collaboration: Rojas et al. (2025) have been modelled via PyAutoLens, which is the standard algorithm for lens modelling in Euclid  [^8] https://github.com/Jammy2211/PyAutoLens (Nightingale & Dye 2015; Nightingale et al. 2018, 2021b). PyAutoLens is an open-source Python package for strong gravitational lensing, which includes both fully automated strong lens modelling of galaxies and galaxy clusters, and tools for simulating samples of strong lenses. In this work, PyAutoLens employs an SIE mass model, consistent with the model used to train LEMON.
[…]
Figure 13: Same as Figs. 8 and 12, but for the Euclid lenses found in the Q1 fields.
Table 4: Same as Table 3, but for the Euclid lenses found in the Q1 fields.
Metrics $R_{\textrm{Ein}}/\textrm{arcsec}$
Bias ($\mu$) 0.05
Lower discr. scatter ($\sigma^{-}$) 0.09
Upper discr. scatter ($\sigma^{+}$) 0.22
RMSE 0.36
MAE 0.20
NMAD 0.14
Coefficient of determination ($R^{2}$) 0.49
7 Speed-up of traditional methods due to LEMON
One possible use of LEMON in the context of strong lensing applications within Euclid, beside fast modelling of strong gravitational lenses, is to speed-up the modelling of strong gravitational lenses by giving traditional lens modelling methods initial starting points from which to start the search.
To perform lens modelling, PyAutoLens adopts the Python library PyAutoFit [^9] https://github.com/rhayes777/PyAutoFit(Nightingale et al. 2021a). The standard modelling pipeline uses nested-sampling algorithms, such as nautilus (Lange 2023), for Bayesian inference of the lens model. To efficiently use the information from LEMON, PyAutoLens can use a gradient ascent optimiser, which fully exploits the information from the initial starting point (Amorisco et al. 2022). The lens light and source galaxy during lens modelling are modelled using a Multi-Gaussian expansion (MGE) technique (see He et al. 2024), independent of any information provided by LEMON, most notably meaning that, when using the gradient ascent optimiser, the centre of the source galaxy uses an initial starting point at (0,0)0″0″(\$\$,\$\$)( 0  ″ , 0  ″ ).
[…]
Finally, to assess whether the optimiser approach needs LEMON in order to correctly model gravitational lenses, we repeated the test by using random starting points for the optimiser. Results show that, with random initialisations, PyAutoLens is unable to successfully recover a correct lens model fit for any of the lenses. Thus, the robust initial starting points provided by LEMON prove to be required in order for PyAutoLens to be sped up by using the gradient ascent optimiser instead of the nested sampling algorithm.
An important benefit of combining LEMON and PyAutoLens lies in their ability to complement each other’s limitations. For instance, the standard PyAutoLens modelling pipeline inferred an incorrect lens model for one system due to misinterpreting additional line-of-sight emission from an unrelated interloper galaxy as part of the lensed source. By using LEMON as a starting point, the initial solution was already aligned with the strong lens, allowing PyAutoLens to refine the model through gradient ascent without being misled by interloper emission. Running independent fits with both the standard PyAutoLens pipeline and the combined LEMON-PyAutoLens approach for the same lens, and selecting the best-fit model from the two therefore provides a viable strategy to minimise failure rates in lens modelling.
8 Conclusions
In this work, we expand the lens modelling machine-learning algorithm LEMON (Gentile et al. 2023) by predicting both the mass and the light of the foreground deflector of Euclid galaxy-galaxy gravitational lenses. The former is modelled as an SIE without external shear, while the latter is a single Sérsic profile. We used mock Euclid lenses with contaminants in the image for both training and testing of the architecture. The main results are as follows.
  • All parameters of the test set are recovered correctly, with very low values of bias and scatter. The parameter that is recovered worst is the Sérsic index, which shows a large scatter and a deviation from the linear trend for high values of the Sérsic index. This suggests that LEMON is able to recover the mass and light parameters for the lens even in the presence of companions contaminating the image.
  • The epistemic uncertainties are systematically lower than the aleatoric uncertainties, which is consistent with a training set that correctly spanned the parameter space.
  • Applied to a selection of Euclidised HST images, LEMON manages to recover Einstein radius, axis ratio, position angle, and lens effective radius of the lens, while we cannot reach any conclusion on the Sérsic index due to low statistics. As expected, in general results are worse than those obtained from the test set, given that Euclidised lenses have a much worse S/N than the simulated lenses, and may not be similar to the simulations used to train the network, as shown in Fig. 10. It should also be noted that both SLACS and ACS lenses have been modelled by assuming a De Vaucouleurs light profile, while LEMON assumes a Sérsic profile with variable values for the Sérsic index, and as such the associated true values for the lens effective radius cannot be compared homogeneously with the LEMON predictions. Moreover, a number of lenses have effective radii bigger than $2\arcsec$, which is the upper limit of the respective training sample for LEMON, and as such their predictions are unreliable extrapolations, and have not been reported. In the future, we will expand the training sample to include lenses with effective radii bigger than $2\arcsec$.
  • Applied to five real galaxy-galaxy Euclid gravitational lenses observed in the Perseus ERO field and the 578 new candidates in the Q1 data (Acevedo Barroso et al. 2024; Euclid Collaboration: Walmsley et al. 2025; Euclid Collaboration: Rojas et al. 2025), for which the value of the Einstein radius is available thanks to classical modelling, we obtain an accurate recovery of the parameter for all lenses.
  • By using the predictions of LEMON as starting points for PyAutoLens, it is possible to speed up the algorithm by up to 26 times compared to the standard Euclid modelling pipeline. This would not be possible without the initial LEMON guesses, because the gradient-ascent optimiser with random initialisations does not produce valid model results for all lenses.
In future works, we plan to further advance the LEMON architecture, by predicting the parameters of more general lens mass models, such as the power-law model. We also plan on enhancing the prediction of certain parameters by deblending the gravitational lens image in lens-only and source-only images (Zhong et al. 2024), and train LEMON separately on the two components. We will also test in more detail the recovery of the lens light parameters, predicting values of effective radius and Sérsic index of real Euclid data, when they will become available. Finally, we will try to upgrade LEMON to also predict the parameters of the source, especially its position, given its usefulness in enhancing the performance of the Euclid standard modelling pipeline.
Q1 data are now public, with DR1 data around the corner. We are therefore preparing to process the enormous amount of upcoming data, and have started applying LEMON to newly discovered strong lenses from the Q1 data release. Results confirm that LEMON is a robust and fast lens modelling tool for Euclid. This will enable the generation of models for the approximately 100 000100000100\,000100 000 expected Euclid lenses in a remarkably short time.
By applying LEMON and PyAutoLens to Euclid gravitational lenses, we will determine the total mass within the Einstein radius, or extrapolated to the effective radius, when lens and source redshifts are available. Combining this with stellar mass estimates from SED fitting (Euclid Collaboration: Enia et al. 2025), we will derive dark matter fractions. We will also provide constraints on the mass density slope and initial mass function. With such a vast statistical lens sample, we will be able to investigate the evolution of these properties in unprecedented detail, as well as the scaling relations which relate them as a function of redshift (Koopmans et al. 2006; Gavazzi et al. 2007; Bolton et al. 2008; Auger et al. 2010; Tortora et al. 2010, 2014, 2018, 2019; Sonnenfeld et al. 2013, 2015, 2017; O’Riordan et al. 2025). By comparing these scaling relations with predictions from cosmological simulations, we will be able to constrain cosmological parameters and physical processes in galaxies (Mukherjee et al. 2018, 2021, 2022; Busillo et al. 2023, 2025; Tortora et al. 2025).
Acknowledgements.
VB and CT acknowledge the INAF grant 2022 LEMON. LL thanks the support from INAF theory Grant 2022: Illuminating Dark Matter using Weak Lensing by Cluster Satellites. This research uses observations made with the NASA/ESA Hubble Space Telescope obtained from the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Inc., under NASA contract NAS 5–26555. This work has made use of the Early Release Observations (ERO) data from the Euclid mission of the European Space Agency (ESA), 2024, https://doi.org/10.57780/esa-qmocze3. This work has made use of the Euclid Quick Release Q1 data from the Euclid mission of the European Space Agency (ESA), 2025, https://doi.org/10.57780/esa-2853f3b. The Euclid Consortium acknowledges the European Space Agency and a number of agencies and institutes that have supported the development of Euclid, in particular the Agenzia Spaziale Italiana, the Austrian Forschungsförderungsgesellschaft funded through BMK, the Belgian Science Policy, the Canadian Euclid Consortium, the Deutsches Zentrum für Luft- und Raumfahrt, the DTU Space and the Niels Bohr Institute in Denmark, the French Centre National d’Etudes Spatiales, the Fundação para a Ciência e a Tecnologia, the Hungarian Academy of Sciences, the Ministerio de Ciencia, Innovación y Universidades, the National Aeronautics and Space Administration, the National Astronomical Observatory of Japan, the Netherlandse Onderzoekschool Voor Astronomie, the Norwegian Space Agency, the Research Council of Finland, the Romanian Space Agency, the State Secretariat for Education, Research, and Innovation (SERI) at the Swiss Space Office (SSO), and the United Kingdom Space Agency. A complete and detailed list is available on the Euclid web site (www.euclid-ec.org).
Appendix A Impact of lack of companions in training set on LEMON performance
Figure 14: Same as Fig. 4, but by using LEMON trained only on non companion lenses and using lenses with companions as the test set.
To further check how the completeness of the training set influences both the predictions and the uncertainties of LEMON, we also trained the algorithm only on the lenses without companions. We split the lenses into 40 0004000040\,00040 000 lenses for the training set, $5000$ lenses for the validation set, and $5000$ lenses for the test set. We then trained LEMON with the same procedure as described in Sect. 4.1, and then tested on a companions test set formed by $5000$ randomly extracted lenses from the 50 0005000050\,00050 000 lenses with companions.
The recovery from the model trained without companions in the training set of the parameters for the companions test set is shown in Fig. 14. From the figure, we can see that the Einstein radius is completely unrecovered, with a very strong bias towards higher predicted radii ($\mu\approx 1\arcsec$) and a very low value of the coefficient of determination, $R^{2}=-14.81$, indicating a poor fit. This suggests that the presence of companions heavily affects the recovery of the Einstein radius, potentially because the contaminants are confused with arc features, or because gravitational effects from the companions alter the lensing characteristics of the images, or both. The axis ratio shows a slightly positive value of $R^{2}$, and a systematic underestimate of the predictions, as seen from the metrics, with $\mu=-0.08$ and $\sigma^{-}=0.14$. This indicates that contaminants have a slight effect on the recovery of the lens shape, likely introducing slight distortions of the image that LEMON was not trained to accommodate. The position angle appears to be well recovered, implying that contaminants have practically no effect on the recovery of this parameter. Although the lens effective radius shows a near-zero mean bias, its $R^{2}$ of $-0.47$ and high upper scatter of $0.21″\sigma^{+}=$$$ highlight that a high scatter induced by the light of the contaminants reduces the model’s ability to predict the effective radius, overestimating systematically the size of the galaxy. The lens Sérsic index shows a very large scatter, with $\textrm{NMAD}=1.07$, reflecting LEMON’s difficulty in recovering the light profile slope, potentially due to additional sources of light in the image that confuse the network. Finally, while the magnitude is the best recovered parameter, with a coefficient of determination $R^{2}=0.77$, it is systematically underestimated ($\mu=-0.29$), indicating that the brightness of contaminants causes LEMON to interpret lenses as brighter than they actually are.
Appendix B Recovery of Einstein mass
As a side test, we trained LEMON to recover the Einstein mass parameter, defined as the total lens mass enclosed within the largest critical curve. This parameter is the effective total mass responsible for the observed lensing effect.
For any axisymmetric lens model, the Einstein mass is related to the (circularised) Einstein radius via the formula
$R_{\textrm{Ein}}=\sqrt{\frac{D_{\textrm{ds}}}{D_{\textrm{s}}D_{\textrm{d}}}\frac{4GM_{\textrm{Ein}}}{c^{2}}}\,,$ (12)
where $D_{\textrm{ds}}$ is the source-lens distance, $D_{\textrm{d}}$ is the distance between the observer and the lens and $D_{\textrm{s}}$ is the distance between the observer and the source. Notice that these distances are angular diameter distances, and thus depend on the redshifts of the source and the lens. We thus expect $M_{\textrm{Ein}}\propto R_{\textrm{Ein}}^{2}$.
The recovery of the Einstein mass for the test set lenses is shown in Fig. 15, with the associated trends for bias and NMAD shown in the lower two panels. From the figure, we can see that the Einstein mass is well recovered over all the lens mass range. It should be noted that this is not expected: given the implicit dependence of Eq. (12) from the redshifts of source and lens, which we did not train the network to recover, the network should have difficulty in recovering the values of Einstein mass through the value of the Einstein radius.
We conjecture that the correct recovery could be due to the fact that the network is learning known or unknown strong correlations between some lensing features and the Einstein mass values, bypassing the need for direct inference of redshift values. Some of these features could be lensing distortion patterns (larger Einstein masses produce more pronounced arcs), size and spread of the lensed images (larger Einstein masses produce wider separations of the lensed images), and brightness of the arcs (larger Einstein masses focus light more intensely, which affects the intensity profiles of the arcs in the image). The network could also be learning implicit correlations between features related to the lens redshift, such as the relative brightness of source and lens, and the redshift of the lens itself. The network could then use these correlations to create a Bayesian mapping that reproduces the relation (12), by learning a distribution of possible mappings between Einstein radius and Einstein mass. Thus, when making predictions, the network sees a particular arc size that often correlates with a higher Einstein mass, and gives a value that peaks around the higher value, with smaller or larger uncertainty associated to how often the value of $R_{\textrm{Ein}}$ is associated to a certain value of $M_{\textrm{Ein}}$ in the training phase.
Figure 16 shows the recovery of the Einstein mass for the SLACS Euclidised lenses sample, which is the only one that has values of Einstein mass reported (Auger et al. 2009, table 4). The recovery of the parameter shows a systematic overestimation, with a mean bias of $\mu=0.23$. This is similar, but slightly larger, to the bias for the recovery of the Einstein radius for the Euclidised lenses.
To further investigate the origin of this systematic effect, in Fig. 17 we show the correlation between the Einstein mass and the Einstein radius. For a fixed redshift, we expect ideally a single power-law curve, corresponding to Eq. (12) with fixed values of $D_{\textrm{ds}}$, $D_{\textrm{s}}$ and $D_{\textrm{d}}$. For different redshifts, in log space we expect a family of curves with fixed slope and varying $y$-intercept, which is what we see in the figure with the training set region. The figure shows that the region of the training set corresponding to the minimum and maximum redshifts of the SLACS sample aligns with the lower boundary of the training set area, where most of the SLACS sample is concentrated. The ERO lenses from Euclid, instead, lie fully within the training-set parameter space, which could explain the better recovery of the Einstein radius for these lenses.
Figure 15: Same as Fig. 4, but for the Einstein mass parameter.
Figure 16: Recovery of the Einstein mass parameter for the Euclidised SLACS lenses, showing how well the predicted parameter from LEMON ($y$ coordinate) reproduces the corresponding true value ($x$ coordinate), obtained from Auger et al. (2009).
Figure 17: Einstein mass as a function of Einstein radius. Grey points are the values for the training set, with black points corresponding to those lenses of the training sample having values of lens redshift within the redshift range of the SLACS sample. Orange squares are the literature parameter values for SLACS lenses, while green triangles are the SLACS predicted values from LEMON. Blue points correspond to the predicted values from LEMON for Euclid ERO lenses considered in Fig. 12.